OpenStudy (anonymous):

Solve 7.5x – 4.5 = 6 – 2.5x.

OpenStudy (anonymous):

x = 0.15 x = 0.3 x = 2.1 x = 1.05 the answers

OpenStudy (uri):

7.5x – 4.5 = 6 – 2.5x. 7.5x+2.5x=6+4.5 10x=10.5 x=10.5/10 x=1.05